2. John has a plan to distribute soft drinks in jimma town for the next year in partnership with Moha
company. He signs a contract with the general manager of Moha soft drinks to operate at a profit
margin of 40 %. Other incremental costs including selling expenses incurred while selling the drinks
in jimma town. John includes these costs at the rate of 20% in addition to the cost of soft drinks.
Further, he incurs an estimated fixed cost of Birr 60,000per year in the operation.
a. What is the linear sales-cost equation
b. What is the breakeven volume of sales in Birr per year
c. What is the retailers profit if he sold soft drinks worth Birr 400,000in a year?
d. If john has sold two crates of soft drinks at Birr 50 each, find the total cost equation
interms of quantities
1
Expert's answer
2020-08-04T18:28:39-0400

As per the question,

Profit margin "=40\%"

Selling expanses "=20\%"

Estimated fixed cost "=60,000" per year

Let the cost price of the soft drink = x,

i) The linear sales cost equation,

"=x+\frac{40x}{100}+\frac{20x}{100}+60000"

"=x+0.4x +0.2x+60000"

"=1.6x+60000"

ii) Breakeven volume of sales in Birr per year "1.6x+60000\/1"

iii) Retailer's profit on the soft drink "=\frac{400000\times 20}{100} =40,000 Rs."